OWG Meeting 02-Jul-2014

From GlueXWiki
Jump to: navigation, search

Location and Time

Room: CC F326-327

Time: 1:30pm-2:30pm


(if problems, call phone in conference room: 757-269-6460)

  1. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 120390084.
  2. To join via a Web Browser, go to the page https://bluejeans.com/120390084.
  3. To join via phone, use one of the following numbers and the Conference ID: 120390084
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. More information on connecting to bluejeans is available.

*note: you'll be muted when first connected to the meeting. press *4 to unmute


  1. Announcements
  2. Major systems reports
    • Trigger
    • DAQ
    • Monitoring
    • Controls
  3. Project Schedule and Activity Prioritization
  4. Mantis Task Tracker

Previous Meeting



David L., Sergey F., Sasha S., Graham H., Carl T., Beni Z., Vardan G., Hovanes E., Sean D., Mark D.


  • Sasha is working on TOF DAQ
    • Able to self-trigger with the CTP
    • Currently reading fADC250 and CAEN TDC
    • Saw cosmic "tracks"
  • BCAL won't be worked on until later next week at the earliest
    • Beni would lim BCAL Trigger for looking at cosmics in CDC and FDC
  • ROCBCAL1 crate had all firmware on FADC250 boards reprogrammed via JTAG (required taking module out)
    • Unsure what happened, but firmware was corrupted on all modules.
    • Caution was urged in the the future since this is a bit of a tedious process


  • Full system tests are progressing
    • Successfully ran with 50 crates, but did see some stalls (most likely related to ROL)
    • Testing will continue every Tuesday and Thursday morning
    • Sasha asked that the 2 TOF crates not be used for the testing on Thursday
  • Still issues with sudden, fatal crashes of DAQ computers
    • Infiniband cards are still prime suspect
  • Currently, there is no means to install OS on new or rebuilt gluon computer that matches others.
    • Some computers have "Desktop" build while others are "Server"
    • Some computers have some additional packages as well.
    • Paul L. will set up a Hall-D "channel" on the satellite server and all gluons will be brought into alignment with that
  • RAID to Tape Library
    • EPICS variable setup that can be alarmed on when available space drops below some threshold
    • System was set up by Paul L. some time ago, but is not yet documented


  • CDC wire positions have been fixed
  • Sergey has made a secondary ROL to create some histograms that can be read via RootSpy. Testing/development still underway


  • FDC channels renamed (Nerses & Vanik)
  • Discussing how to handle discriminator scalers with Hall-B

Action Items